OpenStudy (anonymous):

What is the slope-intercept form of the equation y – 7 = –5/2(x + 4) What is the slope-intercept form of the equation y – 7 = –5/2(x + 4) @Mathematics

jimthompson5910 (jim_thompson5910):

y – 7 = –5/2(x + 4) y – 7 = –5/2x -5/2(4) y – 7 = –5/2x -20/2 y - 7 = -5/2x - 10 y = -5/2x - 10+7 y = -5/2x - 3

OpenStudy (anonymous):